# Natasha Kirova    Serguei Brazovskii

### v1

@@ 1 Introduction
Now we turn to the conceptually more complicated case where the density of excitons and the order parameter $\rho$ are of the same origin; they are additive and statically indistinguishable. This dualism is illustrated by physics of phase transitions from a (quasi)neutral to ionic states in donor-acceptor molecular stacks (see [37] for the latest review and the analysis of the phase diagram, [6, 38] for a review of PIPT experiments, [39] for their later development, and our relevant discussion in [40]). In these compounds, the high-T state already shows a charge alternation $\rho_{1}$ among donor and acceptor sites which is natural because of their different affinities. But at lower T the phase transition takes place to the ”ionic” state characterized by a higher $\rho_{2}>\rho_{1}$ (in units of the electron charge $e$, typically $\rho_{2}\approx 0.7,\rho_{1}\approx 0.3$). Apparently, this transition is not symmetry breaking, but commonly also another, symmetry breaking, order parameter enters the game. This is a spontaneous dimerization of bonds $u$ (akin to (spin)Peierls effect) on top of the either built in or spontaneously augmented dimerization of sites; we shall return to this issue below in Section 4.
@@
The charge transfer present in the ground state can be viewed as a condensate of excitations composed with e-h pairs which otherwise need to be excited across the excitation gap. The appropriate wave function of the e-h state is composed with all possible eigenstates including their free conducting states and bound ones - the Wannier-Mott excitons. But the fortunate fact, that, in the material of the primary interest, the e-h continuum gap lies three times higher than the energy $E_{ex}$ of the excitons, allows us to concentrate only on the later ones. Thus, we arrive at the picture of the Excitonic Insulator suggested long time ago for more conventional semiconductors. The incoherent optical pumping adds a density of excitons which might be additive to the one frozen in to the ground state. But if the added excitons have a time to be cooled down to the BEC state, then both the BE and the EI condensates are described by the unique wave function $\Psi$. This integrity is demonstrated by the duality of theoretical descriptions of the BEC of excitons and of the EI [41]. The formulations differ only by the monitoring parameters: the mean density for the BEC versus the chemical potential for the EI. Microscopically the inerpolation among the BEC and the EI was considered in [42, 43].
@@
Since we concentrate upon a situation with the first order transition, then in equilibrium, without pumping, the energy functional $W(\rho)$ must have two minima: the stable one at $\rho_{1}$ and the metastable one at $\rho_{2}$ separated by the barrier $W_{b}=W(\rho_{b})$ at $\rho_{b}$. The first order phase transition happens at the temperature $T_{c}$ when $W(\rho_{1})=W(\rho_{2})$.  In spite of the essential jump in $\rho$ observed experimentally, there are clear indications that the first order phase transition is close to a second order one. This proximity helpfully allows to use the Landau expansion for the system energy functional which (for a homogeneous state) can be written as
@@
The density $n_{inc}$ of the incoherent cloud of higher energy excitons is taken into account here for further purposes to which we shall return in Section 4. The energy $W_{\rho}$ depends only on the exciton density; its derivative gives the thermodynamic definition of the exciton energy
$\frac{\partial W}{\partial\rho}=U(\rho)=E_{ex}^{0}+a\rho+b\rho^{2}$ (8)
The term $W_{phase}(\Psi)$ is a special energy contribution depending on the phase of $\Psi$ which we shall discuss just below. Notice in (7) the linear dependence of the energy upon the explicit order parameter $\rho$. Actually the energy starts quadratically in therms of the hidden parameter $\Psi$. Also, being small, $\rho$ cannot turn negative because $\rho=\Psi^{\ast}\Psi>0$ which constraint cannot be seen thermodynamically.
Figure 1: Four regions for $W(\rho)$ and $U(\rho)$. I an II: decreasing $U(\rho)>0$, attraction of excitons. II and III: $U(\rho)<0$, spontaneous creation of excitons. III and IV: increasing $U(\rho)>0$, repulsion of excitons
The energy $W(\rho)$ can have two minima: the stable one which we define as $\rho_{1}=0$ and a metastable one at $\rho=\rho_{2}$ separated by the barrier at $\rho_{b}$. This standard form is significant in terms of the excitons as showing four different regimes shown in Fig.1. In the region I $W(\rho)$ grows with $\rho$, hence the exciton energy $E_{ex}(\rho)>0$, but it decreases with increasing $\rho$ which corresponds to the attraction of excitons. In the regions II,III $E_{ex}\eqslantless 0$ indicating the regime of the EI state. Region II and III are separated by the inflation point $\rho^{\ast}$of $W(\rho)$ where $E_{ex}<0$ is minimal; at $\rho>\rho^{\ast}$, the Eex, being still negative, starts to grow which indicates on the excitons’ repulsion. In the region IV we pass the metastable point where $E_{ex}(\rho_{2})=0$ and it grows with increasing of $\rho$ with the feature of the BEC.
Integration of the BEC of excitons and the EI approaches allows for the combined description for the dynamics of excitons and the evolution of the phase transition after the pumping. It include both thermodynamic and dynamic effects of the PIPT. Still, adopting the concept of the EI we arrive at the enigma. The dualism of the excitonic density $|\Psi|^{2}$ and the thermodynamic order parameter $\rho$ do not seem to be consistent. Within the thermodynamic picture, the order parameter $\rho$ is a single real non-conserved field. The expected equation allows for unrestricted evolution towards the energy minimum in $\rho$:
@@
The charge transfer density from excitons $n_{ex}=|\Psi|^{2}$ is also the real field which seems to be equivalent or additive to $\rho$. Nevertheless, its evolution is described by a complex wave function $\Psi=|\Psi|\exp(i\zeta)$ of the BEC. The corresponding evolution of the excitonic condensate contains the hidden degree of freedom, the phase $\zeta$. The evolution of $\zeta$ is measurable through the instantaneous energy of the exciton (defined dynamically as $E_{ex}=-\hbar\partial_{t}\zeta$) which in the homogeneous regime (or in average over the space) coincides with the thermodynamic definition $E_{ex}\Rightarrow U={dW}/{dq}$.
@@
To make the evolution of the fields $\Psi$ and $\rho$ compatible, we need to take into account the processes which do not conserve the number of particles, i.e. given by phase-dependent interactions. They appear from anomalous matrix element of Coulomb interactions transferring two electrons across the gap, from filled to empty band [41] and have been studied intensively for bi-layer systems ([44] with a recent return of interests in [45]). That means the simultaneous creation or annihilation of two e-h pairs, i.e. creation/destruction of two excitons from/to the vacuum. Being virtual usually (the energy is not conserved), these transition amplitudes acquire nonzero averages for macroscopic concentrations in the BEC or EI states.
Figure 2: Normal and anomalous Coulomb interactions
Finally, the excitonic insulator free energy acquires an additional contribution which fixes the phase, hence violates the particles number conservation.
$\Delta W_{phase}(\Psi)=\frac{1}{2}(A^{\ast}\Psi^{2}+A\Psi^{\ast 2})=A\rho\cos[2(\zeta-\alpha)]$ (10)
@@
By the definition of the EI, the phase fixing terms are small, $A\ll 1$, hence the total complex order parameter $\Psi$ still needs to be exploited. The relative value of anomalous $A$ terms, with respect to the main Coulomb energy $E_{b}$, is of the order of $A/E_{b}\propto(a/R)^{d}$ where $a$ is the lattice spacing, $R\gg a$ is the exciton radius, and $d$ is the space dimension which is small by definition for Wannier–Mott excitons. For donor-acceptor chains, the bare value of $A$ may not be very small as $R$ is only of a few $a$ and $d=1$. But the fact that the system is nearly one–dimensional results in strong phase fluctuations which reduce functions periodic in $\zeta$. Then the effective value $A\rightarrow A\langle exp(2i\zeta)\rangle\approx A\exp(-2\langle\zeta^{2}\rangle)$ can be small and even be renormalized to zero (in the context of excitons in two-dimensional structures, effects of the spacial phase incoherence where discussed in [11]).
At a low but still macroscopic density of residual excitons, $\rho\rightarrow 0$, $U\rightarrow E_{ex}^{0}$, the energy of a single excitons is shifted down to $E_{A}=\sqrt{(E_{ex}^{0})^{2}-|A|^{2}}$. The wave function oscillates now with two mirror frequencies: ±$\pm\omega_{A}=\pm E_{A}/\hbar$ where the energy $E_{A}<E_{ex}^{0}$ is shifted down with respect to the bare $E_{ex}^{0}$. The eigenfunction of the single exciton appears as the interference:
@@ 3.2 Evolution equations for the generic excitonic insulator
We shall adopt the Gross–Pitaevskii type equation which is applicable when, for all relevant states of bosons, their occupation numbers are much bigger than unity. For excitons on a $d$–dimensional lattice, the condition that their mean density per lattice site is small is $n_{ex}\gg(T/D)^{d}$. With the exciton bandwidth $D\sim 10^{2}\ K$, this inequality can always be satisfied for a typical experimental temperatures $T\sim 10^{1}$. Even if the initial value of $n_{ex}$ is not sufficiently high, the kinetics of cooling feeds the low energy states, then the Gross–Pitaevskii theory becomes applicable sooner or later. This is an advantage of the fast PIPT technique, where the time of observations is shorter than the recombination life time of excitons.
@@
Here appear two terms breaking the conservation of total number of excitons: the nondissipative term $\varpropto A$ and the common dissipative one $\sim\Gamma$. The relaxation rate $\Gamma$ has a sophisticated behavior: depending on $\rho$, $\Gamma$ evolves from a constant $\Gamma=1/\tau_{rec}$ at vanishing $\rho$ (the single-particle recombination of rare excitons) to $\Gamma\propto\rho$ at moderate $\rho$ (Bose–Einstein statistics and also Auger processes, well proved in semiconductors). Nevertheless, around the second minimum of $W$ at $(\rho\approx\rho_{2})$ of the high density equilibrium phase, we expect that $\Gamma(\rho)\rightarrow 0$ since there is no channel for decay. (We neglect the evaporation over the barrier towards the $\rho=0$.) We can interpolate among these regimes writing $\Gamma=G(\rho)\rho U(\rho)/\hbar$, where $G(\rho)$ is some structureless dimensionless function of $\rho$ (which we shall choose as a constant for the numerical modeling). In the region II,III of Fig.1 $\Gamma<0$ corresponding to the creation of excitons - the amplification of the number of excitons instead of their decay. Another way to reason is to notice that the energy relaxation terminates when $\partial_{t}\zeta=0$ which leads to another form of $\Gamma$ equivalent in a space homogeneous regime:
@@ 3.3 Space homogeneous regime or zero dimension
$\partial_{t}\rho$ $=-\Gamma\rho+2\left|A\right|\rho\sin(2\zeta)=G\rho^{2}\partial_{t}\zeta+\left|A\right|\rho\sin(2\zeta)$ (15)
$=-(G/\hbar)\rho^{2}(U(\rho)-|A|\cos(2\zeta))+2\left|A\right|\rho\sin(2\zeta).$
The static solution $\partial_{t}\Psi=0$ becomes possible taking into account the effect of the phase locking, when
@@
where $W(\rho)$ is normalized to have the bare exciton energy $E_{ex}^{0}=U(0)=1$. With these parameters, we are below the thermodynamic phase transition to the excitonic insulator state but it can exist as a metastable state (the minimum of $W(\rho)$ at $\rho_{2}\approx 1$).
@@
Figure 3 shows time dependencies of the phase and the amplitude, and the trajectory as a parametric polar plot for $\rho(\zeta)$. The sub-barrier pumping leaves the system in the unlocked regime and it relaxes to the virgin ground state with $\rho\rightarrow 0$ and $E_{ex}(t)\rightarrow E_{ex}^{0}$. In the super-barrier regime, $\zeta$ and $\rho$ grow reaching the locking transition where the exciton energy $E_{ex}(t)\rightarrow E_{loc}$ gradually freezes at the thermodynamic equilibrium while the phase is locked via the dynamical transition at an allowed value $\pi(n+1/2)$. For both cases the curves are superimposed by oscillations originated via the $A$ term by destruction/creation of pairs of excitons.
Figure 3: Time evolution of $\rho(t)$, $\zeta(t)$, $\partial_{t}\zeta(t)$ and the polar trajectory $\rho(\zeta)$
3.4 Inhomogeneous regimes and space-time patterns.
@@
Now we open a possibility for a space-inhomogeneous solution considering a one-dimensional system with a length $L$, $-L/2<x<L/2$. The dynamical interplay of various fields results in transient effects like the self–trapping of excitons and the consequent nucleation of regions with increased exciton density. The effect of self-trapping of excitons results in the possibility of local phase transformations already above a threshold $\rho^{\ast}<\rho_{b}$ which is lower than the barrier threshold $\rho_{b}$ for the homogeneous regime. The reason is the collective self–trapping [47] of excitons similar to self–focusing in optics. For $\rho_{ini}<\rho^{\ast}$ the evolution of $\rho$ and $\Psi$ still shows a transient self-trapping, which nevertheless levels out at the long time. For the intermediate initial pumping $\rho^{\ast}<\rho_{ini}<\rho_{b}$, the self-trapping results in a dynamical formation of domains with different states of $\Psi$. Domains grow with time, they experience reflections from the boundaries, collisions and merging. That is accompanied by formation and collapse of domain and domain walls and appearance of propagating wave fronts. The numerical results for a sub-barrier pumping are presented at Fig.4.
Figure 4: Plots for $\rho(t,x)$ at different levels of pumping. A: $\rho_{ini}<\rho_{b}<\rho^{\ast}$. B: $\rho_{b}<\rho_{ini}<\rho^{\ast}$.

### v2

@@ 3.1 Generic Excitonic Insulator
Now we turn to the conceptually more complicated case where the density of excitons and the order parameter, with a magnitude $\rho$, are of the same origin; they are additive and statically indistinguishable. This dualism is illustrated by physics of phase transitions from a (quasi)neutral to ionic states in donor-acceptor molecular stacks (see [36] for the latest review and the analysis of the phase diagram, [6, 37] for a review of PIPT experiments, [38] for their later development, and our relevant discussion in [39]). In these compounds, the high-T state already shows a charge alternation $\rho_{1}$ among donor and acceptor sites which is natural because of their different affinities. But at lower T the phase transition takes place to the “ionic” state characterized by a higher $\rho_{2}>\rho_{1}$ (in units of the electron charge $e$, typically $\rho_{2}\approx 0.7,\rho_{1}\approx 0.3$). Apparently, this transition is not symmetry breaking, but commonly also another, symmetry breaking, order parameter enters the game. This is a spontaneous dimerization of bonds $u$ (akin to (spin)Peierls effect) on top of the either built in or spontaneously augmented dimerization of sites; we shall return to this issue below in Section 4.
@@
The charge transfer present in the ground state can be viewed as a condensate of excitations composed with e-h pairs which otherwise need to be excited across the excitation gap. The appropriate wave function of the e-h state is composed with all possible eigenstates including their free conducting states and bound ones—the Wannier-Mott excitons (called equivalently the “charge-transfer excitons”). But the fortunate fact, that, in the material of the primary interest, the e-h continuum gap lies three times higher than the energy $E_{ex}$ of the excitons, allows us to concentrate only on the later ones. Thus, we arrive at the picture of the Excitonic Insulator suggested long time ago for more conventional semiconductors. The incoherent optical pumping firstly gives rise to a gas of uncorrelated excitons which density might be additive to the one frozen in to the ground state. But if the initial excitons have a time to be cooled down to the BEC state, then both the BE and the EI condensates are described by the unique wave function $\Psi$. This integrity is demonstrated by the duality of theoretical descriptions of the BEC of excitons and of the EI [40]. The formulations differ only by the monitoring parameters: the mean density for the BEC versus the chemical potential for the EI. Microscopically, the interpolation among the BEC and the EI was considered in [41, 42].
@@
Since we concentrate upon a situation with the first order transition, then in equilibrium, without pumping, the energy functional $W(\rho)$ must have two minima: the stable one at $\rho_{1}$ and the metastable one at $\rho_{2}$ separated by the barrier $W_{b}=W(\rho_{b})$ at $\rho_{b}$. The first order phase transition happens at the temperature $T_{c}$ when $W(\rho_{1})=W(\rho_{2})$. In spite of the essential jump in $\rho$ observed experimentally, there are clear indications that the first order phase transition is close to a second order one. This proximity helpfully allows to use the Landau expansion for the energy functional which (for a homogeneous state) can be written as
@@
In (6 the density $n_{incoh}$ of the incoherent cloud of higher energy excitons is taken into account for further purposes to which we shall return in Section 4. The term $W_{phase}(\Psi)$ in (5) is a special energy contribution depending on the phase of $\Psi$ which we shall discuss in this section below. The energy $W_{\rho}$ depends only on the exciton density; its derivative gives the thermodynamic definition of the exciton energy
$\frac{\partial W_{\rho}}{\partial\rho}=U(\rho)=E_{ex}^{0}+a\rho+b\rho^{2}$ (7)
The term $W_{phase}(\Psi)$ in (5) is a special energy contribution depending on the phase of $\Psi$ which we shall discuss just below. Notice in (6) the linear dependence of the energy upon the explicit order parameter $\rho$. Actually the energy starts quadratically in therms of the hidden parameter $\Psi$. Also, being small, $\rho$ cannot turn negative because $\rho=\Psi^{\ast}\Psi>0$ which constraint cannot be seen thermodynamically.
The energy $W(\rho)$ can have two minima: the stable one which we have defined in (6) as $\rho_{1}=0$ and a metastable one at $\rho=\rho_{2}$ separated by the barrier at $\rho_{b}$. This standard form is significant in terms of the excitons as showing four different regimes indicated in Figure 1. In the region I $W(\rho)$ rows with $\rho$, hence the exciton energy $E_{ex}(\rho)>0$, but it decreases with increasing $\rho$ which corresponds to the attraction of excitons. In the regions II,III $E_{ex}\eqslantless 0$ indicating the regime of the EI state. Region II and III are separated by the inflection pnt $\rho^{\ast}$ of the function $W(\rho)$ where $E_{ex}<0$ is minimal; at $\rho>\rho^{\ast}$, the Eex, being still negative, starts to grow which indicates on the excitons’ repulsion. Region II and IV are separated by the metastability point where $E_{ex}(\rho_{2})=0$ and it keeps growing with increasing of $\rho$ with the feature of the BEC.
Figure 1: Plots of the energy $W(\rho)$ (red) and its derivative – the potential $U(\rho)$. Four regions of different stability are separated by vertical dashed lines indicating the critical values $\rho_{b},\rho^{*},\rho_{2}$. I and II: decreasing $U(\rho)>0$ – the attraction of excitons, II and III: $U(\rho)<0$ – spontaneous creation of excitons, III and IV: increasing $U(\rho)>0$ – repulsion of excitons.
Integration of the BEC of excitons and the EI approaches allows for the combined description for the dynamics of excitons and the evolution of the phase transition after the pumping. It includes oth thermodynamic and dynamic effects of the PIPT. Still, adopting the concept of the EI we arrive at the enigma. The dualism of the excitonic density $|\Psi|^{2}$ and the thermodynamic order parameter $\rho$ do not seem to be consistent. Within the thermodynamic picture, the order parameter $\rho$ is a single real non-conserved field. The expected equation (recall the conventional Equation (3)) allows for unrestricted evolution towards the energy minimum in $\rho$:
@@
The charge transfer density from excitons $n_{ex}=|\Psi|^{2}$ is also the real field which seems to be equivalent or additive to $\rho$. Nevertheless, its evolution is described by the complex wave function $\Psi=|\Psi|\exp(i\zeta)$ of the BEC. The corresponding evolution of the excitonic condensate contains the hidden degree of freedom, the phase $\zeta$. The evolution of $\zeta$ is measurable through the instantaneous energy of the exciton (defined dynamically as $E_{ex}=-\hbar\partial_{t}\zeta$) which in the homogeneous regime (or in average over the space) coincides with the thermodynamic definition $E_{ex}\Rightarrow U={dW}/{d\rho}$.
@@
To make the evolution of the fields $\Psi$ and $\rho$ compatible, we need to take into account the processes which do not conserve the number of particles, i.e., given by phase-dependent interactions. They appear from anomalous matrix element of Coulomb interactions transferring two electrons across the gap, from filled to empty band [40] which have been studied intensively for bi-layer systems ([43] with a recent return of interests in [44]). That means the simultaneous creation or annihilation of two e-h pairs, i.e., creation/destruction of two excitons from/to the vacuum. The contributions of responsible interband Coulomb interaction, together with the principle diagonal one, are shown in Figure 2. Being virtual usually (the energy is not conserved), these transition amplitudes acquire nonzero averages for macroscopic concentrations in the BEC or EI states.
Figure 2: Contributions of normal (at the left diagram) and anomalous (at the right one) Coulomb interactions to the energy of the EI. The dashed line shows the e-h interaction. The semi-circular lines with arrows correspond to anomalous Green functions $\propto\Psi,\Psi^{*}$ which mix electronic (e) and hole (h) creation/annihilation operators.
Finally, the EI free energy acquires an additional contribution which fixes the phase, hence violates the particles number conservation.
$\Delta W_{phase}(\Psi)=\frac{1}{2}(A^{\ast}\Psi^{2}+A\Psi^{\ast 2})=|A|\rho\cos[2(\zeta-\alpha)]$ (9)
@@
By the definition of the EI, the phase fixing terms are small, $A\ll 1$, hence the total complex order parameter $\Psi$ still needs to be exploited. The relative value of anomalous $A$ terms, with respect to the main Coulomb energy $E_{b}$, is of the order of $A/E_{b}\propto(a/R)^{d}$ (where $a$ is the lattice spacing, $R\gg a$ is the exciton radius, and $d$ is the space dimension) which is small by definition for Wannier–Mott excitons. For donor-acceptor chains, the bare value of $A$ may not be very small as $R$ is only of a few $a$ and $d=1$. But the fact that the system is nearly one–dimensional results in strong phase fluctuations which reduce functions periodic in $\zeta$. Then the effective value $A\rightarrow A\langle exp(2i\zeta)\rangle\approx A\exp(-2\langle\zeta^{2}\rangle)$ can be small and even be renormalized to zero (in the context of excitons in two-dimensional structures, effects of the spacial phase incoherence where discussed in [11]).
At a low but still macroscopic density of residual excitons, $\rho\rightarrow 0$, $U\rightarrow E_{ex}^{0}$, the wave function oscillates now with two mirror frequencies: ±$\pm\omega_{A}=\pm E_{A}/\hbar$ where the energy $E_{A}<E_{ex}^{0}$ is shifted down with respect to the bare $E_{ex}^{0}$. The eigenfunction of the single exciton appears as the interference:
@@ 3.2 Evolution Equations for the Generic Excitonic Insulator
We shall adopt the Gross–Pitaevskii type equation which is applicable when, for all relevant states of bosons, their occupation numbers are much bigger than unity. For excitons on a $d$–dimensional lattice, this condition reads that their mean density per lattice site is $n_{ex}\gg(T/D)^{d}$. With the exciton bandwidth $D\sim$$10^{2}\ K$, this inequality can be satisfied for a typical experimental temperatures $T\sim 10^{1}$. Even if the initial value of $\rho$ is not sufficiently high, the kinetics of cooling feeds the low energy states, then the Gross–Pitaevskii theory becomes applicable sooner or later. This is an advantage of the fast PIPT technique, where the time of observations is shorter than the recombination life time of excitons.
@@
[…]
@@ 3.3 Space Homogeneous Regime or Zero Dimension
$\partial_{t}\rho$ $=-\Gamma\rho+\left|A\right|\rho\sin(2\zeta)=G\rho^{2}\partial_{t}\zeta+\left|A\right|\rho\sin(2\zeta)$
$=-(G/\hbar)\rho^{2}(U(\rho)-|A|\cos(2\zeta))+\left|A\right|\rho\sin(2\zeta).$ (14)
The static solution $\partial_{t}\Psi=0$ becomes possible taking into account the phase locking effect when
@@
where $W(\rho)$ is normalized to have the bare exciton energy $E_{ex}^{0}=U(0)=1$. With these parameters, we are below the thermodynamic phase transition to the EI state but it can exist as a metastable state (the minimum of $W(\rho)$ at $\rho_{2}\approx 1$).
@@
Figure 3 shows time dependencies of the amplitude, the phase, the exciton energy, and also the trajectory for $\rho(\zeta)$. The sub-barrier pumping leaves the system in the unlocked regime and it relaxes to the virgin ground state with $\rho\rightarrow 0$ and $E_{ex}(t)\rightarrow E_{ex}^{0}$ with linearly growing $-\zeta$. In the super-barrier regime, $\rho$ grows reaching the locking transition where the exciton energy $E_{ex}(t)\rightarrow E_{loc}$ gradually freezes at the thermodynamic equilibrium while the phase is locked via the dynamical transition at an allowed value $\zeta=\pi(n+1/2)$. For both cases the curves are superimposed by oscillations originated via the $A$ term by destruction/creation of pairs of excitons. The calculations presented in Figure 3 have been actually performed within a more general scheme taking into account a distributed pumping (the green line in the plots); we shall return to these figures in Sec.5.2.
Figure 3: Upper row: subcritical pumping, lower row: Supercritical pumping. Left panels: plots for the excitons density $\rho(t)$ – red, the excitons energy $E_{ex}(t)$ – black, the phase $\zeta(t)/\pi$ – blue, and the distributed pumping profile $n(t)$ – green. Right panels: Parametric polar plots $(\rho,\zeta)$.
3.4 Inhomogeneous Regimes and Space-Time Patterns
@@
Now we open a possibility for space-inhomogeneous solutions of Equations (11) and (12) considering one- and two-dimensional systems. The dynamical interplay of various fields gives rise to transient effects like the self–trapping of excitons and the consequent nucleation of regions with increased exciton density. The effect of self-trapping of excitons results in the possibility of local phase transformations above a threshold $\rho_{st}<\rho_{b}$ which is still lower than the barrier threshold $\rho_{b}$ for the homogeneous regime. The reason is the collective self–trapping [46] of excitons similar to self–focusing in optics. For the intermediate initial pumping $\rho_{st}<\rho_{ini}<\rho_{b}$, the self-trapping results in a dynamical formation of domains with different states of $\Psi$. Domains grow with time, they experience reflections from the boundaries, collisions and merging. That is accompanied by formation and collapse of domain and domain walls and appearance of propagating wave fronts. The numerical results for a sub-barrier pumping are presented in Figure 4. Its left panel shows an example for a 1D regime $\rho(x,t)$; this picture accompanies numerous illustrations which have been already presented in [35]. The central and the left panels show results for a 2D system modeled upon a circle with homogeneous initial conditions. The central panel shows the three-domain profile $\rho(x,y,t_{last})$ taken as a snapshot at the last moment $t_{last}$ of the modeled evolution. The right panel shows this evolution in the form of the time-sliced density plot of $\rho(x,y,t_{n})$ at different moments $t_{n}$. We see a roots-like picture where many domains are formed initially to experience a sequence of mergings till the stationary three-domains state.
Figure 4: Plots for an inhomogeneous evolution. Left panel: 1D regime $\rho(x,t)$; central panel: the profile $\rho(x,y,t_{last})$ taken as a snapshot at the last moment $t_{last}=1000$ of the modeled evolution; right panel: the time-sliced density plot of $\rho(x,y,t_{n})$ at different moments $t_{n}$.
Future modeling for higher space dimensions will be of a certain principle interest in view of contradictory expectations in theory of dynamical solitons in dimensions $D\geq 3$. On one side, the Derrick theorem claims an instability of stationary solitons; on another side the NLSE shows stable solitons as oscillatory solutions [47]. The last case is closer to our model in view of intrinsically oscillating nature of the EI with allowed pairs creation.